OpenStudy (anonymous):

find the inverse for y=x+7÷2

OpenStudy (anonymous):

WELCOME TO OPENSTUDY! The inverse is the opposite. So for this I think it would be y=x-7X2

OpenStudy (campbell_st):

if the question is \[y = \frac{ x+ 7}{2}\] then swap x and y \[x =\frac{y + 7}{2}\] now make y the subject